OpenStudy (anonymous):

A cookie factory uses 1/6 of a bag of flour in each batch of cookies. The factory used 1/2 of a bag of flour yesterday. How many batches of cookies did the factory make? A. 1/12 batch B. 3 batches C. 12 batches D. 2 batches

OpenStudy (anonymous):

@SamsungFanBoy @SolomonZelman @welshfella

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

@pooja195

OpenStudy (mrnood):

every batch takes 1/6 of a bag how many times must you use 1/6 until you have totalled 1/2?

OpenStudy (anonymous):

A i think

OpenStudy (mrnood):

you thing 1/12 of a batch? is 1/2 MORE than 1/6? Every time they make a batch they use 1/6 bag. keep adding 1/6 until you get to a total of 1/2 how many times do they use 1/6 to get to 1/2 try writing 1/2 = ??/6

OpenStudy (anonymous):

3

OpenStudy (anonymous):

is it 3 @MrNood

OpenStudy (welshfella):

yes 1/2 = 3/6 1/6 + 1/6 + 1/6 = 3/6

OpenStudy (anonymous):

@welshfella that isnt one of the choices so now idk

OpenStudy (welshfella):

I know I was just illustrating how you might work it out There are 3 1/6's in 1/3 therefore the answer is 3

OpenStudy (welshfella):

* 3 1/6's in 1/2 i should have said.

OpenStudy (anonymous):

thank you lots!